OpenStudy (anonymous):

Harriet had a 2 1/4-foot length of board. She will use 1 2/3 feet for a project today. How many feet of board will remain

OpenStudy (anonymous):

2 1/4 = 9/4 and 1 2/3 = 5/3 board remaining = 9/4 - 5/3 = 27/12 - 20/12 = 7/12

OpenStudy (anonymous):

9/4 - 5/3 = 7/12

OpenStudy (anonymous):

\[\left(2\frac{1}{4}\right)-\left(1\frac{2}{3}\right)=\frac{7}{12} \]